Circular No. 8767

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2006 U4 (SHOEMAKER-LEVY) R. H. McNaught and D. M. Burton report the recovery of comet

P/1991 V1 = 1991b_{1} = 1991 XVIII on images taken with the 0.5-m Uppsala Schmidt telescope at Siding Spring, the comet’s images on Oct. 26.5 UT being fairly weak, with a coma of FWHM 6′′ and an ill-defined faint tail 10′′ long in p.a. 120^{o}. At the request of the Central Bureau, the comet was observed on Oct. 27.4 by A. C. Gilmore at Mt. John Observatory (0.6-m f/6.4 reflector), who in poor seeing described the comet as a diffuse spot of FWHM 4′′; the final position is of poor quality, with the image located between star trails. The indicated correction to the prediction by S. Nakano (MPC 48384 and Comet Handbook 2006) is ΔT = +8.0 days.

2006 UT  α_{2000}  δ_{2000}  Mag. Observer
Oct. 26.46792 19^{h}23^{m}14.^{s}36 −39^{o}35′06′′.7 18.2 McNaught
26.51716 19 23 25.63 −39 34 08.1 17.4  _{”}
27.40930 19 26 52.56 −39 16 40.2  _{”}
27.41528 19 26 53.93 −39 16 32.6  _{”}
27.43316 19 26 57.96 −39 16 06.8 18.7 Gilmore
27.44152 19 26 59.92 −39 15 57.4 18.6  _{”}
27.44991 19 27 01.81 −39 15 46.7 18.7  _{”}
27.47842 19 27 08.58 −39 15 16.1  _{”}

The following orbital elements linking the two apparitions are by B. G. Marsden:

Epoch = 1991 Oct. 31.0 TT

T = 1991 Oct. 13.8642 TT  ω = 333.^{o}1337
e = 0.705912  Ω = 37.9308 }2000.0
q = 1.132412 AU  i = 16.8569

a = 3.850584 AU n^{o} = 0.1304410 P = 7.556 years

Epoch = 2006 Dec. 11.0 TT

T = 2006 Nov. 25.0013 TT  ω = 333.^{o}5580
e = 0.706643  Ω = 37.8728 }2000.0
q = 1.127551 AU  i = 16.9267

a = 3.843617 AU n^{o} = 0.1307959 P = 7.535 years

2006 October 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green